OpenStudy (anonymous):

1/3 divided by 5/6 =

OpenStudy (anonymous):

\[\frac{ \frac{ a }{ b } }{ \frac{ c }{ d } } = \frac{ a }{ b } * \frac{ d }{ c } \] Try this and you will get your answer

OpenStudy (anonymous):

\[\frac{ 1 }{ 3 } * \frac{ 6 }{ 5 }\]

OpenStudy (anonymous):